A 10000kg rail road car travelling at a speed of 24.0m/s strikes an identical car at rest. If the cars look together as a result of the collision, what is their common speed immediately after collision?
1
Expert's answer
2021-03-14T19:19:25-0400

Let's apply the law of conservation of momentum:


"m_1v_1=(m_1+m_2)v_c,""v_c=\dfrac{m_1v_1}{m_1+m_2}=\dfrac{10000\ kg\cdot24\ \dfrac{m}{s}}{10000\ kg+10000\ kg}=12\ \dfrac{m}{s}."